OpenStudy (anonymous):

What number must you add to complete the square? x^2+20x=33

OpenStudy (inkyvoyd):

Do you know how to complete the square?

OpenStudy (inkyvoyd):

HINT x^2+2*10*x+( )=33+( )

OpenStudy (anonymous):

Nope, what is the formula?

OpenStudy (inkyvoyd):

(x+a)^2=x^2+2ax+a^2

OpenStudy (inkyvoyd):

you want to write the left hand as this, and the right hand should be completely constant.

OpenStudy (inkyvoyd):

(can be a number, but no x or x^2)

OpenStudy (apoorvk):

hmm... so there is '20x'. so fo it to be a the '2ab' thingy in the expansion of "(a+b)^2 = a^2 +b^2 +2ab", 20x = 2*x*20. a=x, so b=20. so b62 = 400 so basically, -33 + *some no.* = 400 now what could that *some no.* be?

OpenStudy (anonymous):

Thanks.

OpenStudy (inkyvoyd):

@apoorvk , nice steal :S

OpenStudy (apoorvk):

Lo;, what steal?